A body starts from rest   moves with a uniform acceleration. The ratio of the distance covered in the nth second to the distançe covered in n second

Grade:12th pass

2 Answers

Vikas TU
7 years ago
For nth second,
x = u + 0.5a(2n – 1)
at rest u = 0  
thus
x = 0.5a(2n-1)            .......(1)
For n second,
y = 0 + 0.5an^2          ........(2)
dividde the eqns,
x/y = (2n-1)/n^2
parth
5 years ago
for nth second
x=u+0.5a(2n-1)
at rest u=0
thus x=0.5a(2n-1)
for n second 
y = 0 + 0.5an^2
ratio=x/y=2n-1/n^2